# Relativistic Roche problem for stars in precessing orbits around a spinning black hole

### 11 Mar 2024 v1

@@ Abstract
Tidal disruptions of stars on the equatorial plane orbiting Kerr black holes have been widely studied. However thus far, there have been fewer studies of stars in inclined precessing orbits around a Kerr black hole. In this paper, by using tensor virial equations, we show the presence of possible resonances in these systems for typical physical parameters of black hole-neutron star binaries in close orbits or of a white dwarf/an ordinary star orbiting a supermassive black hole. This suggests the presence of a new instability before the tidal disruption limit is encountered in such systems.
@@ I Introduction
The classical Roche problem studies the equilibrium and stability of a self-gravitating fluid in a binary. Its main result is the presence of a critical orbital distance, the so-called Roche limit, below which there is no stable configuration, i.e., mass shedding and tidal disruption of the star are induced [1, 2]. Extensions of these results to eccentric orbits have also been the focus of some studies [3, 4]. Furthermore, the concept of the Roche limit has been extended to general relativistic scenarios, such as when the binary companion is a black hole (BH) [5, 6, 7, 8]. The previous studies for this focused on the star in circular equatorial orbits around the BH, and to our knowledge, the only exception is [9].
In this paper, we study the case of a star in circular orbits on the non-equatorial plane around a spinning BH. To study this, we assume that the star is incompressible and utilize the tensor virial formalism developed by Chandrasekhar [2]. This formalism has generally been used to study the assumed ellipsoidal shape of a self-gravitating fluid and provided a number of knowledge on the equilibrium and stability of the stars.
@@
This paper is organized as follows: In Sec. II, we present the working assumptions and the tensor virial equations adapted to our system, by which we derive the equilibrium solution of a star on an equatorial orbit around a BH. In Sec. III, we present our method for the perturbative expansion, which we apply up to the second order in $\varepsilon$. In Sec. IV, we show the presence of new resonances in these systems, using the equations derived in Sec. III. Finally, we discuss the implications of our findings in Sec. V. Throughout this paper, we use the geometrical units of $c=G=1$ where $c$ and $G$ are the speed of light and gravitational constant, respectively.
@@ II.1 System
To study the tidal effect on the stars by the tidal force from the BH, we prepare Fermi-normal coordinates on the geodesics [10]. The tidal force to the star by the BH is evaluated on the local inertial frame on the Fermi-normal coordinates. This allows us to treat the problem in the similar way to Newtonian hydrodynamics.
@@
In the inertial frame defined above, we refer to the axes of the corresponding orthonormal basis as 1, 2, and 3. We assume that the axis 3 points in the same direction as the orbital angular momentum of the fluid or in other words, the direction of $(\partial/\partial\theta)^{\mu}$. Note that our 2 and 3 axes agree with the 3 and 2 axes of [10], and that we changed the direction of our 2-axis. Assuming that the stellar radius, $R_{\mathrm{star}}$, is much smaller than the curvature radius of the BH, $r_{0}$, one can write the tidal tensor as [10]
@@ II.2 Tensor virial equations
As already stated in Sec. I, we employ the tensor virial equations [16, 2] to study the equilibrium state of stars orbiting a Kerr BH. The hydrodynamics equation for the fluid on the local inertial frame is written as
@@
where $X_{i}=(X_{1},X_{2},X_{3})$ denote the coordinates in the local inertial frame, $U_{i}$ is the velocity field of the star, $P$ the pressure, and $\phi$ the Newtonian potential of the fluid, which obeys the Poisson equation as
@@
where $x_{i}$ and $u_{i}$ denote the coordinates and the velocity field of the star in the rotating frame, $\delta_{ij}$ is the Kronecker delta, $\epsilon_{ijk}$ is the completely antisymmetric tensor in 3 dimension, and $\dot{\Omega}=d\Omega/d\tau$.
@@
Equation (34) is interpreted as the evolution equation of the angular momentum. Equation (32) will serve as our basis for the exploration of the influence of the tidal force on stars orbiting a Kerr BH.
@@ II.3 Equatorial plane case
When the star orbits a BH on its equatorial plane, the equations become analytically solvable. Although this case was already studied in [8], we will here restate the main results, as they are the zeroth-order solutions for the main problem in this paper.
@@
In this paper, we pay attention to the cases of $f=0$ and 1. For $f=0$, the fluid has no internal motion in the rotating frame. In the inertial frame, the fluid’s internal motion comes only from the rotation. On the other hand, for $f=1$, the fluid has a certain vorticity in the rotating frame, while in the inertial frame, the fluid does not have any circulation. We refer to this case as irrotational Roche-Riemann ellipsoid (IRRE) [2]. It has been shown that this velocity configuration is typically (approximately) satisfied for binary neutron stars and BH-NS binaries in the late inspiraling stage [17, 18].
In the case of the equatorial orbits, Eq. (35) becomes
@@
After fixing $f$ and eliminating $p_{c}/(\pi\rho^{2}a_{1}^{2})$, we can determine $\alpha_{2}$ and $\tilde{\Omega}$ as functions of $\alpha_{3}$. We plot the the relations between $\tilde{\Omega}$ and $\alpha_{3}$ for $f=0$ and $1$ in Fig. 1, which agree with those in [8].
@@ III.1 Method for the expansion
We consider a spherical orbit with always the same radius $r_{0}$ but different values of $\theta$. Following [14], to study the orbit of the star, we fix $r_{0}$ and $\cal{C}$ and deduce the other parameters from ${\cal{R}}=0=d{\cal{R}}/dr$ at $r=r_{0}$. Thus, $E$, $L$, and $K$ depend on the inclination angle. These changes occur at the second order in $\varepsilon_{0}$ (see appendix).
@@
The components only contain even-order contributions, if its indices are even on the component 3, and odd-order contributions, if its indices are odd on the component 3.
@@
We do the same for the displacement coefficients $\xi_{ij}$. By considering the symmetry $x_{3}\to-x_{3}$, one can constrain the expression of the displacement. We thus have that if the indices of the displacement are odd on the component 3, then the series expansion only contains terms of odd power in $\varepsilon$, and likewise for indices even on component 3, the series expansion only contains terms of even power in $\varepsilon$.
@@ III.2 First-order calculations
$\frac{d^{2}}{d\tau^{2}}V_{i;j}-2\sum_{k=1}^{3}Q_{jk}\frac{dV_{i;k}}{d\tau}=-\sum_{l=1}^{3}((Q^{2})_{jl}V_{i;l}+(Q^{2})_{il}V_{j;l})-2\pi\rho B_{ij}V_{ij}+\pi\rho a_{i}^{2}\delta_{ij}\sum_{l=1}^{3}A_{il}V_{ll}\\ +\Omega_{0}^{2}V_{ij}-\delta_{i3}\Omega_{0}^{2}V_{3j}+\delta_{ij}\delta\Pi+2\sum_{l=1}^{3}\epsilon_{il3}\Omega_{0}\left(\frac{d}{d\tau}V_{l;j}+\sum_{k=1}^{3}(Q_{lk}V_{j;k}-Q_{jk}V_{l;k})\right)-\sum_{k=1}^{3}\tilde{C}^{(0)}_{ik}V_{kj}-\sum_{k=1}^{3}\tilde{C}^{(1)}_{ik}I_{kj}.$ (64)
As before, we can write the corresponding symmetrized and antisymmetrized equations. We only need to consider 4 of these equations at first order. Indeed only the equations for the indices $(i,j)=(1,3),(3,1),(2,3),(3,2)$ contain the non-zero contribution from the inertia tensor.
@@ III.3 Second-order calculations
To find the second-order equations, we need to take into account second-order corrections to $\Omega$ and $\dot{\Omega}$ for the equations. We denote the former as $\Omega=\Omega_{0}+\Omega^{(2)}+O(\varepsilon_{0}^{4})$. The second-order corrections to $\Omega$ and $\dot{\Omega}$ are given in the appendix.
For the second-order perturbation, we only need to consider the even terms on the component 3. We then have 5 equations. By adding Eq. (49), we obtain 6 equations for 6 variables, $V_{1;1}$, $V_{2;2}$, $V_{3;3}$, $V_{1;2}$, $V_{2;1}$, and $\delta\Pi$.
@@
For the complete expression of these terms, we refer the leader to the appendix.
@@
The differential equations for the constant term must be considered with more care. If we proceed as before, we obtain again linear equations $\sum_{k=1}^{3}M_{ik}^{(2,0)}V_{k}^{(2,0)}=C_{i}^{(2,0)}$. The matrix of this equation $M_{ik}^{(2,0)}$ is the same as $M_{ik}^{(2,2)}$ when we set $\omega_{\theta}=0$. However, this matrix is never inversible as its third and fourth line are colinear.
However, in this case the term in the right hand side, $C_{i}^{(2,0)}$, is also simpler because $C_{3}^{(2,0)}=C_{4}^{(2,0)}=0$ (see Eq. (88)). This would imply that we have 4 equations for 5 variables. We have thus a certain freedom for setting $V_{1;2}^{(2,0)}$ and $V_{2;1}^{(2,0)}$. We could set them both to 0.
@@ IV.1 Determination of resonances
However, during our research, we found that these matrices could have determinant 0 for particular equilibrium states. Furthermore, the terms in the right hand side were not in the image of the matrices. Thus the linear equations were not solvable for the particular cases in which the determinant vanishes. This would indicate that the ansatz used in these cases are inappropriate. One has then a resonance. To solve the equations, a better ansatz would be: $\vec{\xi}=(\vec{\xi}_{0}+\vec{\xi}_{1}\tau)e^{i\omega_{\theta}\tau}$. In this case, the displacements would eventually not be infinitesimal. Thus our perturbative expansion breaks down at such points.
@@
In our present context, the displacement associated with an oscillatory mode gets excited by the variation of the $\theta$-parameter of the orbit. It will then grow linearly and eventually, it will not be infinitesimal anymore. Figures 3 and 4 show the displacement vectors associated with the excited modes, assuming $a/m=0.8$. Again the spin parameter does not have an important impact on the qualitative aspect of the resonant displacement.
@@
The resonance in the first- and second-order displacements is always encountered slightly outside the Roche/Roche-Riemann limits. For BH-NS binaries in precessing orbits, thus, the resonant deformation may be excited just prior to the tidal disruption and the tidal disruption may be assisted by the prior resonance if the relativistic effect on the self-gravity of the star, which we do not take into account in the present study, does not qualitatively change the property of the resonance.
@@ V Discussion
One interpretation of our results is that as an orbiting star approaches gradually a Kerr BH, e.g., by the radiation reaction of gravitational waves, it will approach one of these resonances before encountering the tidal disruption limit. The closer it gets to it, the more its displacement associated with it will grow. At some point before the resonance, the displacement will not be infinitesimal anymore. Our equations are then not valid.
@@
To our knowledge, these results have never been reported before. A major counterpoint to our analysis is its apparent reliance on some questionable assumptions, like the incompressibility hypothesis. However, as the general mechanism for the resonance holds true, we expect resonances even with more realistic equations of state. Indeed, whatever the equation of state is, we have characteristic modes for our star. We may then observe a resonance between these modes and the oscillation present in the tidal tensor. Thus, while the equations can change, the resulting phenomenon should not qualitatively. Of course, this explanation is just tentative and should be confirmed by a more serious analysis of the question.
@@ Acknowledgements.
We thank the member of Computational Relativistic Astrophysics group at the Max Planck Institute for Gravitational Physics for discussion. This work was in part supported by Grant-in-Aid for Scientific Research (grant Nos. 20H00158 and 23H04900) of Japanese MEXT/JSPS.
Appendix A Formula for second-order equations in $\varepsilon$

### 14 Mar 2024 v2

@@ Abstract
Tidal disruptions of stars on the equatorial plane orbiting Kerr black holes have been widely studied. However thus far, there have been fewer studies of stars in inclined precessing orbits around a Kerr black hole. In this paper, we use the tensor virial equations to show the presence of possible resonances in these systems for typical physical parameters of black hole-neutron star binaries in close orbits or of a white dwarf/an ordinary star orbiting a supermassive black hole. This suggests the presence of a new instability before the tidal disruption limit is encountered in such systems.
@@ I Introduction
The classical Roche problem studies the equilibrium and stability of a self-gravitating fluid in a binary system. Its main result is the presence of a critical orbital distance, the so-called Roche limit, below which there is no stable configuration, i.e., mass shedding and tidal disruption of the star are induced [1, 2]. Extensions of these results to eccentric orbits have also been the focus of some studies [3, 4]. Furthermore, the concept of the Roche limit has been extended to general relativistic scenarios, such as when the binary companion is a black hole (BH) [5, 6, 7, 8]. The previous studies for this focused on the star in circular equatorial orbits around the BH, and to our knowledge, the only exception is [9].
In this paper, we study the case of a star in a circular orbit on the non-equatorial plane around a spinning BH. To study this, we assume that the star is incompressible and utilize the tensor virial formalism developed by Chandrasekhar [2]. This formalism has generally been used to study the assumed ellipsoidal shape of a self-gravitating fluid and provided a number of knowledge on the equilibrium and stability of the stars.
@@
This paper is organized as follows: In Sec. II, we present the working assumptions and the tensor virial equations adapted to our system, by which we derive the equilibrium solution of a star on an equatorial orbit around a BH. In Sec. III, we present our method for the perturbative expansion, which we apply up to the second order in $\varepsilon$. In Sec. IV, we show the presence of new resonances in these systems, using the equations derived in Sec. III. Finally, we discuss the implications of our findings in Sec. V. Throughout this paper, we use the geometrical units of $c=G=1$ where $c$ and $G$ are the speed of light and gravitational constant, respectively. We will also not use the Einstein summation convention.
@@ II.1 System
To study the effect on the stars of the tidal force from the BH, we prepare Fermi-normal coordinates on the geodesics [10]. The tidal force acting on the star by the BH is evaluated in the local inertial frame on the Fermi-normal coordinates. This allows us to treat the problem in the similar way to Newtonian hydrodynamics.
@@
In the inertial frame defined above, we refer to the axes of the corresponding orthonormal basis as 1, 2, and 3. We assume that the axis 3 points in the same direction as the orbital angular momentum of the fluid or in other words, the direction of $(\partial/\partial\theta)^{\mu}$. Note that our 2 and 3 axes agree with the 3 and 2 axes of [10], and that we changed the direction of our 2-axis. Assuming that the stellar radius, $R_{\mathrm{star}}$, is much smaller than the orbital separation to the BH, $r_{0}$, one can write the tidal tensor as [10]
@@ II.2 Tensor virial equations
As already stated in Sec. I, we employ the tensor virial equations [16, 2] to study the equilibrium state of stars orbiting a Kerr BH. The hydrodynamic equation for the fluid on the local inertial frame is written as
@@
where $X_{i}=(X_{1},X_{2},X_{3})$ denote the coordinates in the local inertial frame, $U_{i}$ is the velocity field of the star, $P$ the pressure, and $\phi$ the Newtonian potential of the fluid. It obeys the Poisson equation as
@@
where $x_{i}$ and $u_{i}$ denote the coordinates and the velocity field of the star in the rotating frame, $\delta_{ij}$ is the Kronecker delta, $\epsilon_{ijk}$ is the completely antisymmetric tensor in 3 dimensions, and $\dot{\Omega}=d\Omega/d\tau$.
@@
Equation (34) is interpreted as the evolution equation of the angular momentum.
Equation (32) will serve as our basis for the exploration of the influence of the tidal force on stars orbiting a Kerr BH.
@@ II.3 Equatorial plane case
When the star orbits a BH on its equatorial plane, the equations become analytically solvable. Although this case was already studied in [8], we will here restate the main results, as they serve as the zeroth-order solutions for the main problem in this paper.
@@
In this paper, we pay attention to the cases of $f=0$ and $f=1$. For $f=0$, the fluid has no internal motion in the rotating frame. In the inertial frame, the fluid’s internal motion comes only from the rotation. On the other hand, for $f=1$, the fluid has a certain vorticity in the rotating frame, while in the inertial frame, the fluid does not have any circulation. We refer to this case as irrotational Roche-Riemann ellipsoid (IRRE) [2]. It has been shown that this velocity configuration is typically (approximately) satisfied for binary neutron stars and BH-NS binaries in the late inspiraling stage [17, 18].
In the case of equatorial orbits, Eq. (35) becomes
@@
After fixing $f$ and eliminating $p_{c}/(\pi\rho^{2}a_{1}^{2})$, we can determine $\alpha_{2}$ and $\tilde{\Omega}$ as functions of $\alpha_{3}$. We plot the the relations between $\tilde{\Omega}$ and $\alpha_{3}$ for $f=0$ and $f=1$ in Fig. 1, which agree with those in [8].
@@ III.1 Method for the expansion
We consider a spherical orbit with always the same radius $r_{0}$ but different values of $\theta$. Following [14], to study the orbit of the star, we fix $r_{0}$ and $\cal{C}$ and deduce the other parameters from ${\cal{R}}=0=d{\cal{R}}/dr$ at $r=r_{0}$. Thus, $E$, $L$, and $K$ depend on the inclination angle $\theta$. These changes occur at the second order in $\varepsilon_{0}$ (see appendix).
@@
The components only contain even-order contributions, if its indices are even in the component 3, and odd-order contributions, if its indices are odd in the component 3.
@@
We do the same for the displacement coefficients $\xi_{ij}$. By considering the symmetry $x_{3}\to-x_{3}$, one can constrain the expression of the displacement. We thus have that if the indices of the displacement are odd in the component 3, then the series expansion only contains terms of odd power in $\varepsilon$, and likewise for indices even in the component 3, the series expansion only contains terms of even power in $\varepsilon$.
@@ III.2 First-order calculations
$\frac{d^{2}}{d\tau^{2}}V_{i;j}-2\sum_{k=1}^{3}Q_{jk}\frac{dV_{i;k}}{d\tau}=-\sum_{l=1}^{3}((Q^{2})_{jl}V_{i;l}+(Q^{2})_{il}V_{j;l})-2\pi\rho B_{ij}V_{ij}+\pi\rho a_{i}^{2}\delta_{ij}\sum_{l=1}^{3}A_{il}V_{ll}\\ +\Omega_{0}^{2}V_{ij}-\delta_{i3}\Omega_{0}^{2}V_{3j}+\delta_{ij}\delta\Pi+2\sum_{l=1}^{3}\epsilon_{il3}\Omega_{0}\left(\frac{d}{d\tau}V_{l;j}+\sum_{k=1}^{3}(Q_{lk}V_{j;k}-Q_{jk}V_{l;k})\right)-\sum_{k=1}^{3}\tilde{C}^{(0)}_{ik}V_{kj}-\sum_{k=1}^{3}\tilde{C}^{(1)}_{ik}I_{kj},$ (64)
with $V_{ij}=V_{i;j}+V_{j;i}$.
As before, we can write the corresponding symmetrized and antisymmetrized equations. We only need to consider four of these equations at first order. Indeed only the equations for the indices $(i,j)=(1,3),(3,1),(2,3),(3,2)$ contain the non-zero contribution from the inertia tensor.
@@ III.3 Second-order calculations
To find the second-order equations, we need to take into account second-order corrections to $\Omega$ and $\dot{\Omega}$ for the equations. We denote the former as $\Omega=\Omega_{0}+\Omega^{(2)}+O(\varepsilon_{0}^{4})$. The second-order corrections to $\Omega$ and $\dot{\Omega}$ are given in the appendix A.
For the second-order perturbation, we only need to consider the even terms on the component 3. We then have five equations. By adding Eq. (49), we obtain six equations for six variables, $V_{1;1}$, $V_{2;2}$, $V_{3;3}$, $V_{1;2}$, $V_{2;1}$, and $\delta\Pi$.
@@
For the complete expression of these terms, we refer the reader to the appendix.
@@
The differential equations for the constant term must be considered with more care. If we proceed as before, we obtain again linear equations $\sum_{k=1}^{3}M_{ik}^{(2,0)}V_{k}^{(2,0)}=C_{i}^{(2,0)}$. The matrix of this equation $M_{ik}^{(2,0)}$ is the same as $M_{ik}^{(2,2)}$ when we set $\omega_{\theta}=0$. However, this matrix is never invertible as its third and fourth line are colinear.
However, in this case the term in the right hand side, $C_{i}^{(2,0)}$, is also simpler because $C_{3}^{(2,0)}=C_{4}^{(2,0)}=0$ (see Eq. (88)). This would imply that we have four equations for five variables. We have thus a certain freedom for setting $V_{1;2}^{(2,0)}$ and $V_{2;1}^{(2,0)}$. We could set them both to 0.
@@ IV.1 Determination of resonances
However, during our research, we found that these matrices could have determinant 0 for particular equilibrium states. Furthermore, the terms in the right hand side were not in the image of the matrices. Thus the linear equations were not solvable for the particular cases in which the determinant vanishes. This would indicate that the ansatz used in these cases are inappropriate.
One has then a resonance. To solve the equations, a better ansatz would be: $\vec{\xi}=(\vec{\xi}_{0}+\vec{\xi}_{1}\tau)e^{i\omega_{\theta}\tau}$. In this case, the displacements would eventually not be infinitesimal. Thus our perturbative expansion breaks down at such points.
@@
In our present context, the displacement associated with an oscillatory mode gets excited by the variation of the $\theta$-coordinate of the orbit. It will then grow linearly and eventually, it will not be infinitesimal anymore. Figures 3 and 4 show the displacement vectors associated with the excited modes, assuming $a/m=0.8$. Again the spin parameter does not have an important impact on the qualitative aspect of the resonant displacement.
@@
The resonance in the first- and second-order displacements is always encountered slightly outside the Roche/Roche-Riemann limits. For BH-NS binaries in precessing orbits, thus, the resonant deformation may be excited just prior to the tidal disruption. The tidal disruption may be assisted by the prior resonance if the relativistic effect on the self-gravity of the star, which we do not take into account in the present study, does not qualitatively change the property of the resonance.
@@ V Discussion
One interpretation of our results is that as an orbiting star approaches a Kerr BH, e.g., by the radiation reaction of gravitational waves, it will approach one of these resonances before encountering the tidal disruption limit. The closer it gets to it, the more the displacement associated with it will grow. At some point before the resonance, the displacement will not be infinitesimal anymore. Our equations are then not valid.
@@
To our knowledge, these results have never been reported before. A major counterpoint to our analysis is its apparent reliance on some questionable assumptions, like the incompressibility hypothesis. However, as the general mechanism for the resonance holds true, we expect resonances even with more realistic equations of state. Indeed, whatever the equation of state is, we have characteristic modes for our star. We may then observe a resonance between these modes and the oscillation present in the tidal tensor. Thus, while the equations can change, the resulting phenomenon should not qualitatively. Of course, this explanation is just tentative and should be confirmed by a more complete analysis of the question.
@@ Acknowledgements.
We thank the member of Computational Relativistic Astrophysics group at the Max Planck Institute for Gravitational Physics for discussions. This work was in part supported by Grant-in-Aid for Scientific Research (grant Nos. 20H00158 and 23H04900) of Japanese MEXT/JSPS.
Appendix A Formula for second-order equations